ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

14. Mathematical Aspects of Computer Science

1991 MS Classification: 5, 68 Kasyanov, Victor, Institute of Informatics Systems, Russia Hierarchical graphs and visual processing Hierarchical graphs as a new graph formalism for representing hierarchical kind of diagramming objects are introduced. Hierarhical graphs are suited for a wide array of applications in computer science. Problems of hierarchical graph drawing are studied. Methods and tools for visual processing in the framework of hierarchical graph models are considered. The work has been partially supported by the RFBR under grant 98-01-748. Reference: V. N. Kasyanov, Hierarchical graphs and graph models: visual processing problems, Problems of Informatics Systems and Programming, Novosibirsk (1998), (In Russion, in print).